# Multi-object Data Integration in the Study of Primary Progressive Aphasia

### v1

@@ 5.2 Competitors and Metrics of Comparison
The simulated data will be used to assess the performance of: (A) identifying influential regions; (B) estimating the true network coefficient ${\boldsymbol{\Theta}}_{t}$ and structural coefficients ${\boldsymbol{\beta}}_{1,t},...,{\boldsymbol{\beta}}_{P,t}$; and (C) quantifying uncertainty in the point estimation of network and structural coefficients. We construct a series of competitors of BMRR to assess (A)-(C) as described below.
@@ 5.2.2 Bayesian competitors
We construct a few Bayesian competitors to compare single modal analysis and joint analysis of modalities without integrating information between objects, with the integrative analysis performed by BMRR. These competitors include:
(a) fitting only the Bayesian network on scalar regression (i.e., the first equation in (3)), referred to as the network response regression (NRR); (b) fitting only the Bayesian structural images on scalar regression (i.e., the second equation in (3), referred to as the structural response regression (SRR); and (c) fitting both of them jointly without integrating them through shared parameters, referred to as the joint regression (JR). For NRR, a region is identified as influential if at least one edge connecting to that region is influential. In SRR, a region is deemed influential if it contains at least one influential voxel. For JR, a region is identified as influential if any voxel in that region or any edge connected to that region is influential.
We implement two additional Bayesian competitors, both capturing joint effects of all network edges and cells of the structural images on the predictor, but do not acknowledge the network topology or the connection between the topology of two sets of objects through the hierarchical constraint. The first Bayesian competitor, referred to as the Spike & Slab, applies an ordinary spike & slab prior (George and McCulloch, 1993) on each $\theta_{p,p^{\prime}}$ and $\beta_{v,p}$. A voxel or an edge will be identified as influential if the posterior probability that the corresponding coefficient equals zero exceeds $0.5$ (Barbieri and Berger, 2004). We also implement a Horseshoe shrinkage prior (Carvalho et al., 2010) on each $\theta_{p,p^{\prime}}$ and $\beta_{v,p}$. Since the Horseshoe prior does not result in exact zeros for the post-burn-in iterates of coefficients, we perform a post-processing approach following Li and Pati (2017) to determine which voxel coefficients and network edges are related to the predictor. For both Spike & Slab and Horseshoe competitors, an ROI $\mathcal{R}_{p}$ is identified as influential if at least a network edge connecting to that ROI or at least on voxel coefficient within the ROI is nonzero.
@@ 5.2.3 Metrics of comparison
We will present the $F_{1}$ score computed from the True Positive Rate (TPR) and the True Negative Rate (TNR) for correctly identifying important regions. The point estimation of every competitor is assessed using mean squared errors (MSE) of estimating the network coefficient ${\boldsymbol{\Theta}}_{t}$ and the structural coefficients ${\boldsymbol{\beta}}_{1,t},...,{\boldsymbol{\beta}}_{P,t}$. Since both the fitted ${\boldsymbol{\Theta}}$ and ${\boldsymbol{\Theta}}_{t}$ are symmetric with zero diagonals, the MSE for both sets of coefficients jointly is given by $[\sum_{p<p^{\prime}}(\theta_{p,p^{\prime},t}-\widehat{\theta}_{p,p^{\prime}})^{2}+\sum_{p=1}^{P}||{\boldsymbol{\beta}}_{p,t}-\widehat{{\boldsymbol{\beta}}}_{p}||^{2}]/(P(P-1)/2+VP)$. The point estimates are taken to be the posterior median for the Bayesian competitors. For Bayesian competitors, we evaluate the length and coverage of 95% credible intervals averaged across coefficients in ${\boldsymbol{\Theta}}$ and present them for all cases. Similar quantities for ${\boldsymbol{\beta}}_{p}$’s are also presented. All results presented over $500$ simulated datasets.
@@ 5.3 Results
Figure 3 presents the $F_{1}$ score for BMRR region identification under each of our simulation scenarios in Table 1. As the voxel density parameter $\upsilon_{t}$ approaches 0, indicating high voxel-sparsity, region identification suffers. Similarly, BMRR’s performance deteriorates with lower SNR. However, for the same $\upsilon_{t}$ and SNR, BMRR performs better in high-dimensional cases than in low-dimensional cases. Additionally, for the same levels of $\upsilon_{t}$, SNR, and $(V,P)$, BMRR performs better with a higher degree of co-localization, since it induces co-localization in its modeling framework. Finally, we observe that the smoothly varying spatial correlation structure over the voxels of the structural images does not significantly impact BMRR’s performance in terms of region identification.
@@
Figure 4 shows boxplots of $F_{1}$ scores for structured Bayesian regressions BMRR, NRR, SRR, and JR under various simulation scenarios. Due to space constraints, we present results for moderate and low SNR values of $1$ and $0.5$, and voxel density parameter $\upsilon_{t}=0.2$ and $0.4$. It is evident that the integrated version, BMRR, consistently outperforms using either or both objects independently. While using both objects in JR improves performance compared to using only one image in NRR or SRR, the improvement is not as significant as when using the integrated BMRR approach. Both BMRR and JR show marginal performance improvements with increasing SNR, increasing co-localization degree, and in high-dimensional settings. In contrast, the improvement is significant for NRR and SRR in the high-dimensional setting. All competitors show better performance with increasing voxel density as well as with decreasing node sparsity within a limit. Additionally, the $F_{1}$ scores do not show any significant perturbation with spatial association between voxels of the structural image.
@@
According to Figure 5, BMRR outperforms every competitor in every scenario in terms of $F_{1}$ score. While Spike & Slab, Horseshoe, and MUA show average $F_{1}$ scores close to 1 as the co-localization degree increases, their scores exhibit wide variability over replications, as demonstrated by the large interquartile ranges. TRR performs poorly across all scenarios. BMRR also outperforms other competitors in terms of estimating network and structural coefficients, as evidenced by the smallest MSE values under all scenarios. Spike & Slab and Horseshoe perform similarly to each other, while MUA and TRR show the poorest performance. Notably, the performance of all competitors remains largely unchanged between spatial and non-spatial scenarios with perfect co-localization. For uncertainty quantification, we observe that all methods offer nominal or slight over-coverage, except for TRR, which shows severe under-coverage. Additionally, BMRR provides narrower 95% credible intervals than its competitors under all scenarios, with the exception of Spike & Slab, which marginally outperforms BMRR in low-dimensional cases.
@@ 6.2 Data analysis results
Figure 6 displays horizontal slices (higher c coordinate refers to the top of the brain) of the GM coefficient map organized by ROI where the colored shading indicates the magnitude and direction of association of the speech rate with z-score GM segmentation probability. Due to the hierarchical nature of the BMRR model, the distribution of influential GM voxels (as captured by their associated coefficients) is determined by the ROI selection and thus reflects the distribution of the selected 46 ROIs discussed above, with concentrations in the SFG, MFG, IFG, IPL, SPL, and BG in a bilateral distribution. The GM coefficient maps provide an additional level of granularity by characterizing the association between the speech rate and GM probability at the voxel level as well as the sparsity of the signal which varies by ROI. Overall, speech rate is positively associated with increases in voxel level GM in the frontal gyrus which affirms that a decrease in GM in these regions is correlated with lower articulation rates (Landin-Romero et al., 2021; García et al., 2022). However, there is heterogeneity in the voxel level signal both within and across ROIs with a substantial number of voxels displaying a negative association of GM with speech rate particularly in the parietal lobe. The effect sizes range from -1.4 to 1.4 across voxels which indicates that a standard deviation shift in speech rate (25 words/minute) can produce a change in z-score GM segmentation probability from between -1.4 and 1.4. These results reveal a dynamic portrait of how GM and neurodegeneration are associated with motor speech and highlight heterogeneity in the direction of the association as well as varying levels of sparsity in influential voxels across ROIs, insights that are missed when data is pooled at the ROI level.
@@
We compare the results from the application of the proposed BMRR to the MUA. Figures 7-8 in the supplementary file show that very few GM voxels and network edges are selected, providing us with very few signals for analysis in MUA. We also contrast BMRR with Horseshoe and JR in terms of the Posterior Predictive Loss Criterion (PPLC) that shows the fit of the three competing models. The results presented in Table 3 show better model fit by BMRR than JR, indicating the advantages of integrating information from objects. Given that the coefficients for the speech rate measure are not very large, BMRR has only a modest advantage over Horseshoe in model fitting statistics. However, Horseshoe proves extremely unreliable in identifying influential ROIs, as it marks all ROIs as influential, rendering the results scientifically uninterpretable. Overall, BMRR develops novel insights on the neuro-degeneration pathway for PPA and outperforms competitors due to accounting for science-driven constraints in the prior structure for coefficients.
@@
In our empirical analysis, we set the hyper-parameter values as $a_{\nu}=b_{\nu}=1=a_{\tau_{\theta}}=b_{\tau_{\theta}}=a_{\tau_{\beta}}=b_{\tau_{\beta}}=1$. To assess the sensitivity to the choice of hyper-parameters, data analysis is conducted with different combinations of hyper-parameters. Specifically, we vary $a_{\nu}=b_{\nu}$ at values $0.1,1,10$. For each such choice, we vary $a_{\tau_{\theta}}=b_{\tau_{\theta}}=a_{\tau_{\beta}}=b_{\tau_{\beta}}$ at values $0.1,1,10$. For each hyper-parameter combination, MSPE, coverage of the 95% predictive interval, and the length of the 95% predictive intervals averaged over all image voxels and network edges are shown. We also compute the Adjusted Rand Index (ARI) between the partitions of influential and uninfluential node indices obtained from a specific combination of hyper-parameters and from the choice $a_{\nu}=b_{\nu}=1$ and $a_{\tau_{\theta}}=b_{\tau_{\theta}}=a_{\tau_{\beta}}=b_{\tau_{\beta}}=1$. All these metrics in Table 4 indicate that BMRR is largely insensitive to the choice of hyper-parameters.
Table 3: Weighted Posterior Predictive Loss Criterion (PPLC) as in as in Gelfand and Ghosh (1998) for values of for BMRR, using the non-integrated version of BMRR (JR) and Horseshoe. $PPLC=G+P$, where G is associated with goodness of fit and P is associated with a penalty related to model complexity.
@@ 7 Conclusion and Future Work
Motivated by multi-modal imaging applications constituting the structural and functional imaging data on patients with PPA, this article develops a regression approach with structural and network-valued objects on a scalar predictor. A novel prior structure is developed jointly on coefficients corresponding to different objects which can simultaneously exploit object characteristics and the linked information between the objects to draw inference on network nodes significantly related to the scalar predictor with uncertainty. The proposed approach is arguably the first statistical multi-modal response regression approach that allows for the flexibility of drawing inference at the ROI level and the voxel level simultaneously and is equipped with identifying brain regions significantly related to the speech rate measuring neuro-degeneration due to PPA with uncertainty. The analysis of PPA data leads to an important understanding of neuro-degeneration patterns associated with PPA.
The study has some important limitations and considerations. Our results leverage the organizing hierarchy of the Brainnetome atlas and its associated ROIs to connect structural and network information. Thus, conclusions might differ had an alternate brain atlas been used. While sensitivity analyses utilizing different brain atlas could be conducted, we emphasize that our proposed technique and its application to the PPA data is based on strong prior neuroanatomical knowledge that the Brainnetome atlas and subsequent parcellation is appropriate for the clinical question of interest. In settings where investigators are unsure of the appropriate atlas, greater emphasis should be placed on assessing the sensitivity of results to the selection of alternate parcellations. In addition, PPA being an infrequently studied disease, our results are based on a modest sample size which requires cautious interpretation of results as discussed in Section 2. As an immediate future work, we will extend our approach to incorporate the spatial correlation in the GM image.
@@ Acknowledgments
Rajarshi Guhaniyogi acknowledges funding from National Science Foundation Grant DMS-2220840 and DMS-2210672; and National Institute of Health Grant 1R01NS131604-01A1. Aaron Scheffler acknowledges funding from the National Science Foundation Grant DMS-2210206 and National Institute of Health Grant 1R01NS131604-01A1. We thank the Memory and Aging Center at the University of California, San Francisco for providing the MR images and the patients and their families for the time and effort they dedicated to the research. The study was supported by Grants from the National Institutes of Health (NINDS R01NS050915, NIDCD K24DC015544, NIA P50AG023501)

### v2

@@ 5.2 Competitors and Metrics of Comparison
The simulated data will be used to assess the performance of: (A) identifying influential regions; (B) estimating the true network coefficient ${\boldsymbol{\Theta}}_{t}$ and structural coefficients ${\boldsymbol{\beta}}_{1,t},...,{\boldsymbol{\beta}}_{P,t}$; and (C) quantifying uncertainty in inference. We construct a series of competitors to assess (A)-(C) as below.
@@ 5.2.2 Bayesian competitors
We construct a few Bayesian competitors to compare single modal analysis and joint analysis of modalities without integrating information between objects, with the integrative analysis performed by BMRR. These competitors include: (a) fitting only the Bayesian network on scalar regression (i.e., the first equation in (3)), referred to as the network response regression (NRR); (b) fitting only the Bayesian structural images on scalar regression (i.e., the second equation in (3), referred to as the structural response regression (SRR); and (c) fitting both of them jointly without integrating them through shared parameters, referred to as the joint regression (JR). For NRR, a region is identified as influential if at least one edge connecting to that region is influential. In SRR, a region is deemed influential if it contains at least one influential voxel. For JR, a region is identified as influential if any voxel in that region or any edge connected to that region is influential.
We implement two additional Bayesian competitors, which do not acknowledge the network topology or the connection between the topology of two sets of objects through the hierarchical constraint. The first Bayesian competitor, referred to as the Spike & Slab, applies an ordinary spike & slab prior (George and McCulloch, 1993) on each $\theta_{p,p^{\prime}}$ and $\beta_{v,p}$. We also implement a Horseshoe shrinkage prior (Carvalho et al., 2010) on each $\theta_{p,p^{\prime}}$ and $\beta_{v,p}$. As in the frequentist competitors, we controlled for False Discovery Rate (FDR). Unlike BMRR, which naturally selects regions, these methods require careful tuning for FDR-based region selection. For the Spike & Slab method, FDR is applied using the indicator variables. However, it is not straightforward for the Horseshoe prior and it requires a rejection region. Since we could not empirically determine this region, we used nested rejection regions (intervals) for individual coefficients, selecting the smallest interval that met the FDR threshold. We control FDR for joint objects similar to the frequentist methods, yielding the best performance.
It is important to highlight that while the Horseshoe prior provides strong theoretical guarantees for estimation and uncertainty quantification of regression coefficients (van der Pas et al., 2017), its theoretical foundation for identifying important predictors remains limited. In particular, the variable selection framework using the Horseshoe prior in our joint regression setting lacks theoretical guarantees for influential region identification with uncertainty quantification. This is a key advantage of BMRR, which not only facilitates region selection but also provides precise uncertainty quantification in this process.
@@ 5.2.3 Metrics of comparison
We will present the $F_{1}$ score computed from the True Positive Rate (TPR) and the True Negative Rate (TNR) for correctly identifying important regions. The point estimation of every competitor is assessed using mean squared errors (MSE) of estimating the network coefficient ${\boldsymbol{\Theta}}_{t}$ and the structural coefficients ${\boldsymbol{\beta}}_{1,t},...,{\boldsymbol{\beta}}_{P,t}$. Since both the fitted ${\boldsymbol{\Theta}}$ and ${\boldsymbol{\Theta}}_{t}$ are symmetric with zero diagonals, the MSE for both sets of coefficients jointly is given by $[\sum_{p<p^{\prime}}(\theta_{p,p^{\prime},t}-\widehat{\theta}_{p,p^{\prime}})^{2}+\sum_{p=1}^{P}||{\boldsymbol{\beta}}_{p,t}-\widehat{{\boldsymbol{\beta}}}_{p}||^{2}]/(P(P-1)/2+VP)$. The point estimates are taken to be the posterior median for the Bayesian competitors. For Bayesian competitors, we evaluate the length and coverage of 95% credible intervals averaged across coefficients in ${\boldsymbol{\Theta}}$ and ${\boldsymbol{\beta}}_{p}$’s. All results presented over $500$ simulated datasets.
@@ 5.3 Results
Figure 3 presents the $F_{1}$ score for BMRR region identification under each of our simulation scenarios in Table 1. As the voxel density parameter $\upsilon_{t}$ approaches 0, indicating high voxel-sparsity, region identification suffers. Similarly, BMRR’s performance deteriorates with lower SNR. However, for the same $\upsilon_{t}$ and SNR, BMRR performs better in high-dimensional cases than in low-dimensional cases. Additionally, for the same levels of $\upsilon_{t}$, SNR, and $(V,P)$, BMRR performs better with a higher degree of co-localization, since it induces co-localization in its modeling framework. Finally, smoothly varying spatial correlation over the voxels of the structural images does not significantly impact BMRR’s performance in terms of region identification.
@@
Figure 4 shows boxplots of $F_{1}$ scores for BMRR, NRR, SRR, and JR. Due to space constraints, we present results for moderate and low SNR values of $1$ and $0.5$, and voxel density parameter $\upsilon_{t}=0.2$ and $0.4$. The integrated version, BMRR, consistently outperforms using either or both objects independently. While using both objects in JR improves performance compared to using only one image in NRR or SRR, the improvement is not as significant as when using the integrated BMRR approach. Both BMRR and JR show marginal performance improvements with increasing SNR, increasing co-localization degree, and in high-dimensional settings. In contrast, the improvement is significant for NRR and SRR in the high-dimensional setting. All competitors show better performance with increasing voxel density as well as with decreasing node sparsity within a limit. Additionally, the $F_{1}$ scores do not show any significant perturbation with spatial association between voxels of the structural image.
@@
According to Figure 5, BMRR outperforms every competitor in every scenario in terms of $F_{1}$ score. While Spike & Slab, Horseshoe, and MUA show average $F_{1}$ scores close to 1 as the co-localization degree increases, MUA scores exhibit wide variability over replications, as demonstrated by the large interquartile ranges. BMRR also outperforms other competitors in terms of estimating network and structural coefficients, as evidenced by the smallest MSE values under all scenarios. Spike & Slab and Horseshoe perform similarly to each other, while MUA and TRR show the poorest performance. Notably, the performance of all competitors remains largely unchanged between spatial and non-spatial scenarios with perfect co-localization. For uncertainty quantification, we observe that all methods offer nominal or slight over-coverage, except for TRR, which shows severe under-coverage. Additionally, BMRR provides narrower 95% credible intervals than its competitors under all scenarios, with the exception of Spike & Slab, which marginally outperforms BMRR in low-dimensional cases.
Notably, our simulation study explores the impact of violating the co-localization assumption on model performance. While we observe a decrease in region identification performance (F1-score), the effects on mean squared error (MSE), coverage, and interval length are less pronounced. Importantly, our method still outperforms other approaches even when co-localization is not strictly met. This demonstrates that while co-localization is a desirable characteristic, it is not a prerequisite for effective performance with our method. These results support our scientific motivation of achieving accurate simultaneous region identification across both modalities, even in the presence of imperfect co-localization.
@@ 6.2 Data analysis results
Figure 6 displays horizontal slices (higher c coordinate refers to the top of the brain) of the GM coefficient map organized by ROI where the colored shading indicates the magnitude and direction of association of the speech rate with z-score GM segmentation probability. Due to the hierarchical nature of the BMRR model, the distribution of influential GM voxels (as captured by the coefficients) is determined by the ROI selection and thus reflects the distribution of the selected 46 ROIs discussed above, with concentrations in the SFG, MFG, IFG, IPL, SPL, and BG in a bilateral distribution. The GM coefficient maps provide an additional level of granularity by characterizing the association between the speech rate and GM probability at the voxel level as well as the sparsity of the signal which varies by ROI. Overall, speech rate is positively associated with increases in voxel level GM in the frontal gyrus which affirms that a decrease in GM in these regions is correlated with lower articulation rates (Landin-Romero et al., 2021; García et al., 2022). However, there is heterogeneity in the voxel level signal both within and across ROIs with a substantial number of voxels displaying a negative association of GM with speech rate particularly in the parietal lobe. The effect sizes range from -1.4 to 1.4 across voxels which indicates that a standard deviation shift in speech rate (25 words/minute) can produce a change in z-score GM segmentation probability from between -1.4 and 1.4. These results reveal a dynamic portrait of how GM and neurodegeneration are associated with motor speech and highlight heterogeneity in the direction of the association as well as varying levels of sparsity in influential voxels across ROIs, insights that are missed when data is pooled at the ROI level.
@@
We compare the results from the application of the proposed BMRR to the MUA. Figures 7-8 in the supplementary file show that very few GM voxels and network edges are selected, providing us with very few signals for analysis in MUA. Table 3 shows better model fit by BMRR than JR in terms of the Posterior Predictive Loss Criterion (PPLC), indicating the advantages of integrating information from objects. Given that the coefficients for the speech rate measure are not very large, BMRR has only a modest advantage over Horseshoe in model fitting statistics. However, Horseshoe proves extremely unreliable in identifying influential ROIs, as it marks all ROIs as influential, rendering the results scientifically uninterpretable.
In our empirical analysis, we set the hyper-parameter values as $a_{\nu}=b_{\nu}=1=a_{\tau_{\theta}}=b_{\tau_{\theta}}=a_{\tau_{\beta}}=b_{\tau_{\beta}}=1$. To assess the sensitivity of hyper-parameters, we vary $a_{\nu}=b_{\nu}$ at values $0.1,1,10$. For each such choice, we vary $a_{\tau_{\theta}}=b_{\tau_{\theta}}=a_{\tau_{\beta}}=b_{\tau_{\beta}}$ at values $0.1,1,10$. For each hyper-parameter combination, MSPE, coverage of the 95% predictive interval, and the length of the 95% predictive intervals averaged over all image voxels and network edges are shown. We also compute the Adjusted Rand Index (ARI) between the partitions of influential and uninfluential node indices obtained from a specific combination of hyper-parameters and from the choice $a_{\nu}=b_{\nu}=1$ and $a_{\tau_{\theta}}=b_{\tau_{\theta}}=a_{\tau_{\beta}}=b_{\tau_{\beta}}=1$. All these metrics in Table 4 indicate that BMRR is largely insensitive to the choice of hyper-parameters.
@@
Table 3: Posterior Predictive Loss Criterion (PPLC) as in as in Gelfand and Ghosh (1998) for values of for BMRR, JR and Horseshoe. $PPLC=G+P$, where G is associated with goodness of fit and P is associated with a penalty related to model complexity.
@@ 7 Conclusion and Future Work
Motivated by multi-modal imaging applications constituting the structural and functional imaging data on patients with PPA, this article develops a regression approach with structural and network-valued objects on a scalar predictor. A novel prior structure is developed jointly on coefficients corresponding to different objects which can simultaneously exploit object characteristics and the linked information between the objects to draw inference on network nodes significantly related to the scalar predictor with uncertainty. The proposed approach is arguably the first statistical multi-modal response regression approach that allows for the flexibility of drawing inference at the ROI level and the voxel level simultaneously and is equipped with identifying brain regions significantly related to the speech rate measuring neuro-degeneration due to PPA with uncertainty. The analysis of PPA data leads to an important understanding of neuro-degeneration patterns for PPA.
The study has some important limitations and considerations. Our results leverage the organizing hierarchy of the Brainnetome atlas and its associated ROIs to connect structural and network information. Thus, conclusions might differ had an alternate brain atlas been used. We emphasize that our proposed technique and its application to the PPA data is based on strong prior neuroanatomical knowledge that the Brainnetome atlas and subsequent parcellation is appropriate for the clinical question of interest. In settings where investigators are unsure of the appropriate atlas, greater emphasis should be placed on assessing the sensitivity of results to the selection of alternate parcellations. In addition, PPA being an infrequently studied disease, our results are based on a modest sample size which requires cautious interpretation of results as discussed in Section 2. As an immediate future work, we will extend our approach to incorporate the spatial correlation in the GM image.
@@ Acknowledgments
Rajarshi Guhaniyogi acknowledges funding from National Science Foundation Grant DMS-2220840 and DMS-2210672; and National Institute of Health Grant 1R01NS131604-01A1. Aaron Scheffler acknowledges funding from the National Science Foundation Grant DMS-2210206 and National Institute of Health Grant 1R01NS131604-01A1. We thank the Memory and Aging Center at the University of California, San Francisco for providing the MR images and the patients and their families for the time and effort they dedicated to the research. The study was supported by Grants from the National Institutes of Health (NINDS R01NS050915, NIDCD K24DC015544, NIA P50AG023501, NIH RF1NS050915)